OpenStudy (anonymous):

If we define a stream function \[\Psi(x,y) = \ln\sqrt{x^{2} + y^2}\] What are the corresponding velocity components of V {vector} = u * i + v * j

OpenStudy (anonymous):

If our stream function is defined in terms of psi, then the u will equal the partial derivative of psi with respect to y and v will equal the partial derivative of psi with respect to x and it's negative. \[\Psi _{y}=(1/\sqrt{x^2+y^2})(1/(2\sqrt{x^2+y^2}))(2y)\] \[\Psi _{y}=y/\sqrt{x^2+y^2}\] The partial of x is identical except the numerator is an x. \[-\Psi _{x}=-x/\sqrt{x^2+y^2}\] v (vector) = \[iy/\sqrt{x^2+y^2}-jx \sqrt{x^2+y^2}\]